If Colleen Mooney invests $5,122.90 now and she will receive $12,900 at the end of 12 years, what annual rate of interest will Colleen earn on her investment? Annual rate of interest %?
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

12,900=5,122.90*(1+r/100)^12

(12,900/5,122.90)^(1/12)=(1+r/100)

(1+r/100)=1.08

r=1.08-1

=8%